OpenStudy (anonymous):

in a collection of nickels and quarters, there are 36 coins. The total value of the coins is $2.90. How many coins of each type are there?

OpenStudy (whpalmer4):

let N be the number nickels and Q be the number of quarters. We have a total of $2.00*100 cents/$1 = 200 cents. The value of the nickels is 5N, and the value of the quarters is 25Q. Can you write the 2 equations that show that we have 12 nickels and quarters, and that the value of those coins is 200 cents?

OpenStudy (anonymous):

x+y=12?

OpenStudy (whpalmer4):

well, we were going to use N for the number of nickels, and Q for the number of quarters, much more descriptive, easier to keep track of what they mean... \[N+Q = 12\]That's a good start. Now you need an equation that relates the number of coins of each type, their values, and the total value (200 cents).

OpenStudy (anonymous):

is x=18 and y=7?

OpenStudy (whpalmer4):

I have no idea what x and y are. I could tell you if you are correct if you tell me what N and Q are, however :-)

OpenStudy (whpalmer4):

Let me ask you this: does 18+7 = 12?

OpenStudy (anonymous):

no

OpenStudy (whpalmer4):

then that can't be a solution...

OpenStudy (whpalmer4):

also, even if you're talking 18 nickels, that's almost $1 right there, and 7 quarters is $1.75...

OpenStudy (jhannybean):

You can buy a bar of ice cream with that much.

OpenStudy (whpalmer4):

Let's get that other equation put together. We have the value of the nickels: 5N We have the value of the quarters: 25Q We have the total value of all the coins: 200 combine those 3 quantities with + or - and =